An object falls from the Transco Tower in Houston and takes 15 seconds to reach the ground. 

A. What is its velocity at impact if air resistance is ignored?


B. How tall is the building? 


C. What is its acceleration at the 2nd second? 


D. What is its acceleration at the 5th second? 


E. If the Transco Tower were actually 3,000 meters tall, how long would an object take to free-fall off of the top of the building? (ignoring air resistance) 




1
Expert's answer
2020-11-05T07:34:00-0500

a)


"v=(9.8)(15)=147\frac{m}{s}"

b)


"h=0.5(9.8)(15)^2=1102.5\ m"

c)


"a=9.8\frac{m}{s^2}"

d)


"a=9.8\frac{m}{s^2}"

e)


"3000=0.5(9.8)t^2\\t=24.74\ s"